# HEST-1k: A Dataset for Spatial Transcriptomics and Histology Image Analysis

### v1

@@ C.1 HEST-1k
Ex-ST Andrusivova and Fan (2023) Brain Visium 5 31,053
Genome-wide spatial expression profiling in formalin-fixed tissues Villacampa et al. (2021) Brain Visium 15 31,053
Human ileum, Visium Mirzazadeh et al. (2021) Bowel Visium 4 33,538
Human squamous cell carcinoma Abalo et al. (2021) Skin Visium 4 33,538
Prostate needle biopsies pre- and post-ADT: Count matrices, histological-, and Androgen receptor immunohistochemistry images Marklund (2022) Prostate Spatial Transcriptomics 24 26,437
Spatially resolved clonal copy number alterations in benign and malignant tissue Erickson et al. (2022) Prostate Visium 23 33,538
Spatially resolved transcriptomic profiling of degraded Mirzazadeh et al. (2023) Bowel Visium 35 17,943
@@
Transcriptome-scale spatial gene expression in the human dorsolateral prefrontal cortexMaynard et al. (2021) Brain Visium 12 33,538
A spatially resolved atlas of the human lung characterizes a gland-associated immune nicheMadissoon et al. (2023) Lung Visium 20 17,922
@@
Tertiary lymphoid structures generate and propagate anti-tumor antibody-producing plasma cells in renal cell cancer Meylan et al. (2022) Lymph node Visium 24 17,943
@@
Charting the Heterogeneity of Colorectal Cancer Consensus Molecular Subtypes using Spatial Transcriptomics: datasets Valdeolivas et al. (2023) Bowel Visium 14 36,601
Demo 10x Visium dataset for STQ zen (2024a) Skin Visium 1 68,886
Spotiphy: generative modeling in single-cell spatial whole transcriptomics zen (2024b) Brain Visium 2 32,285
@@ C.2 HEST-Library
The HEST-Library helps transform unstructured spatial transcriptomics and histology data into a unified format. An overview of the HEST-Library is provided in Suppl. Figure 3. An example of fiducial detection is presented in Suppl. Figure 4.
Figure 3: Overview of HEST-Library functionalities. HEST was designed to transform legacy data scrapped in multiple public repositories, such as NCBI, into unified and actionable objects that can easily be integrated into multimodal computational pipelines.
Figure 4: Fiducial detection and automatic alignment in Visium samples. Corner fiducials on 6.5$\times$6.5mm and 11mm×$\times$11mm Visium slides are automatically detected with a finetuned Yolov8 model, the spots coordinates are then derived if at least 3 of the 4 corner fiducials are detected. This process enables automatically estimating the pixel resolution.
@@ C.3 HEST-Benchmark
Gene selection, Random Forest, and Ridge regression models: We learn a regression model that maps the patch embeddings of each encoder to its corresponding gene expression profile. The Random Forest model is based on 70 trees (sklearn implementation). The Ridge regression uses a fixed $L2$ regularization coefficient $\lambda$ to $100/MC$, where $M$ is the embedding dimension and $C=50$ is the number of targets trained with the Regularized Least-Squares Routine solver (sklearn implementation). We train the regression models to predict a panel of 50 most variable genes tailored to each task. Specifically, for each task, we select the 50 most variable genes across all spots and samples, after excluding the genes that have non-zero counts in less than 10% of the spots.
@@
Task 1 Janesick et al. (2023) IDC Breast 4 4 Xenium
@@
Task 5 Valdeolivas et al. (2023) COAD Colon 3 6 Visium
Task 6 Valdeolivas et al. (2023) READ Rectum 2 4 Visium
Task 7 Meylan et al. (2022) ccRCC Kidney 24 24 Visium
Task 8 Giraud et al. (2022) HCC Liver 2 2 Visium
Task 9 LUAD Lung 2 2 Xenium
Task 10 Liu et al. (2022) IDC Axillary lymph nodes 4 4 Visium
@@
Task 1: Expression prediction in invasive ductal carcinoma (breast cancer, IDC). We used all publicly available Xenium samples available on 10x Genomics (“FFPE Human Breast using the Entire Sample Area”, 2 patients) and two samples published in  Janesick et al. (2023) (2 patients). All samples are FFPE sections imaged with the Xenium pipeline v1.
@@
Task 2: Expression prediction in prostate adenocarcinoma (prostate cancer, PRAD). We used all 23 Visium samples (fresh frozen sections) from 2 patients published in Erickson et al. (2022). Both patients were diagnosed with prostatic acinar adenocarcinoma with a (4+3) Gleason score (ISUP group 4).
@@
Task 3: Expression prediction in pancreatic adenocarcinoma (pancreatic cancer, PAAD). We used 3 samples from 3 different patients from 10x Genomics (“FFPE Human Pancreas with Xenium Multimodal Cell Segmentation” and “Pancreatic Cancer with Xenium Human Multi-Tissue and Cancer Panel”). All samples are FFPE sections processed with Xenium pipeline v1.
@@
Task 4: Expression prediction in skin cutaneous melanoma (skin cancer, SKCM). We used 2 samples from 2 different patients from 10x Genomics website (“Human Skin Data with Xenium Human Multi-Tissue and Cancer Panel”). All samples are FFPE sections processed with Xenium pipeline v1.
@@
Task 5: Expression prediction in colon adenocarcinoma (colon cancer, COAD). We used 6 COAD samples from 3 different patients published in Valdeolivas et al. (2023). All samples are fresh frozen sections processed with Visium.
@@
Task 6: Expression prediction in rectal adenocarcinoma (rectum cancer, READ). We used 4 READ samples from different 2 patients published in Valdeolivas et al. (2023). All samples are fresh frozen sections processed with Visium.
@@
Task 7: Expression prediction in clear cell renal cell carcinoma (kidney cancer, ccRCC). We used all 24 ccRCC samples from 24 different patients published in Meylan et al. (2022). All samples are fresh frozen sections processed with Visium.
Task 8: Expression prediction in hepatocellular carcinoma (liver cancer, HCC). We used 2 HCC liver samples from 2 different patients published in Giraud et al. (2022). All samples are fresh frozen sections processed with Visium.
@@
Task 9: Expression prediction in lung adenocarcinoma (lung cancer, LUAD). We used 2 LUAD samples from 2 different patients from 10x genomics (“Preview Data: FFPE Human Lung Cancer with Xenium Multimodal Cell Segmentation”). All samples are fresh frozen sections processed with Xenium pipeline v1.
@@
Task 10: Expression prediction in axillary lymph nodes in IDC patients. We used 4 axillary lymph node samples from 2 IDC patients published in Liu et al. (2022). All samples are fresh frozen sections processed with Visium.
@@
Table A12: State-of-the-art foundation models for histology evaluated on HEST-Benchmark. ResNet-50: 23M parameters, ResNet-18: 11M, ResNet-152: 795M, ViT-B: Base (86M), L: Large (307M), G: Giant (1.13B)
Name Number of Number of Magnification Model Training
slides patches recipe
ResNet50-IN Lu et al. (2021b) N/A 1M N/A ResNet-50 Supervised
KimiaNet Riasatian et al. (2021) 7K 240K 20x ResNet-18 Supervised
Ciga Ciga et al. (2022) UNK UNK 10x, 20x ResNet-50 Supervised
CTransPath Wang et al. (2021) 32k 17M 10x UNK MoCov2
Remedis Azizi et al. (2023) 10 10M 20x ResNet-152 iBOT
Phikon Filiot et al. (2023) 6k 43.3M 20x ViT-B iBOT
PLIP Huang et al. (2023) N/A 800k Many ViT-B CLIP
UNI Chen et al. (2024) 100k 100M 20x ViT-L DINOv2
CONCH Lu et al. (2024) N/A 1.17M Many ViT-B CoCa
GigaPath Xu et al. (2024) 171k 1.3B 20x ViT-G CoCa
@@
We provide a brief description of each patch encoder assessed on the HEST-Benchmark.
@@
ResNet50 (IN) Lu et al. (2020): This model uses a ResNet50 backbone He et al. (2016) trained on ImageNet dataset Deng et al. (2009). Following prior work Lu et al. (2020), the patch embedding corresponds to the deep representation at the penultimate layer before classification.
Ciga Ciga et al. (2022): This model uses a ResNet-50 backbone He et al. (2016) trained on public histology datasets such as TCGA and Cameleyon16 using SimCLR.
@@
CTransPath Wang et al. (2022): This model uses a “Tiny” Swin Transformer backbone Liu et al. (2021) with a window size of 14 (Swin-T/14, 28 million parameters) pretrained on TCGA and PAIP datasets (17 million images) using MoCoV3 Chen et al. (2021).
@@
Remedis Azizi et al. (2023): This model uses a ResNet-152$\times$2 (232 million parameters) initialized with the “Big Transfer”-medium protocol Kolesnikov et al. (2020) on ImageNet-22K and pretrained with SimCLR Chen et al. (2020) on TCGA.
@@
Phikon Filiot et al. (2023): This model uses a Vision Transformer-Base (ViT-B, 86 million parameters) Dosovitskiy et al. (2021) trained on TCGA data using iBOT Zhou et al. (2022).
PLIP Huang et al. (2023): This model uses a ViT-B (86 million parameters) trained on 208 thousand pathology images paired with natural language descriptions extracted from Twitter and trained using CLIP Radford et al. (2021).
@@
UNI Chen et al. (2024): This model uses a ViT-Large (ViT-L, 307 million parameters) Dosovitskiy et al. (2021) trained on 100 million histology images (over 100,000 slides) from proprietary and public data using DINOv2 Oquab et al. (2023).
@@
CONCH Lu et al. (2024): This model uses a ViT-B (86 million parameters) trained on 1.17 million histology image–caption pairs extracted from online educational and research resources using CoCa Yu et al. (2022).
@@
GigaPath Xu et al. (2024): This model uses a ViT-Giant (1.13 billion parameters) trained on 1.3 billion image patches from 171,189 WSIs at 20$\times$ magnification using DINOv2.
@@ Appendix D HEST for multimodal representation learning
CONCH-FT model, a ViT-Base model initialized with CONCH weights, was fine-tuned for 50 epochs using a cosine learning rate scheduler, with a base learning rate of $10^{-4}$ for the image encoder and $10^{-3}$ for the expression encoder. Only the last 3 layers of the model were fine-tuned, with a layer-wise learning decay rate of 0.7. For training with the infoNCE loss, a contrastive temperature of $10^{-2}$ and batch size of 1,024 pairs of patch and transcriptomics were used. A combination of random horizontal/vertical flip and color jittering was employed for image patch augmentation.
The rank of the embedding space, also referred to as smooth rank measure Garrido et al. (2023), measures the quality of the embeddings produced from encoders trained in unsupervised or self-supervised manners. Given the patch embedding matrix $H\in\mathcal{R}^{N\times d}$ and $d<N$, where $N$ is the number of patches and $d$ is the feature dimension, we compute the rank as the entropy of the $d$ L1-normalized singular values of $H$.
@@ Appendix E HEST for discovery
Cells were segmented and classified using CellViT Hörst et al. (2024). To find the gene expression profile of each neoplastic cell, we matched each cell to its corresponding cell index in Xenium by assigning the index for which the distance between the cell centroids was the smallest. After matching all neoplastic cells, only those cells for which the assignment was unique were kept. After this filtering step, an average of 91% of the cells per sample were kept while 9% of the cells were discarded. The discarded cells likely correspond to over-segmentation errors by CellViT.
@@ F.1 Motivation for dataset creation
  • Why was the dataset created? HEST-1k was designed with three key applications: (1) multimodal representation learning of histology and transcriptomics, (2) biomarker discovery, and (3) benchmarking of histopathology patch encoder models. Despite many publicly available resources, no existing unified and user-friendly formatting was available to bring ST in the world of deep learning.
  • What (other) tasks could the dataset be used for? Are there obvious tasks for which it should not be used? Users are welcome to propose new creative ways to use the dataset. Users are not allowed to try to retrieve patient information from the existing data. A dedicated section is allocated to discuss ethical considerations and intended usage.
  • Has the dataset been used for any tasks already? If so, where are the results so others can compare (e.g., links to published papers)? All samples that were made public as part of a publication (peer-review or not) are reported in the metadata attached with HEST-1k.
  • Who funded the creation of the dataset? The work was funded by an R35 grant.

### v2

@@ C.1 HEST-1k
Ex-ST [andrusivova2023exst] Brain Visium 5 31,053
Genome-wide spatial expression profiling in formalin-fixed tissues [villacampa2021genome] Brain Visium 15 31,053
Human ileum, Visium [mirzazadeh2021human] Bowel Visium 4 33,538
Human squamous cell carcinoma [abalo2021human] Skin Visium 4 33,538
Prostate needle biopsies pre- and post-ADT: Count matrices, histological-, and Androgen receptor immunohistochemistry images [marklund2022prostate] Prostate Spatial Transcriptomics 24 26,437
Spatially resolved clonal copy number alterations in benign and malignant tissue [erickson2022spatially] Prostate Visium 23 33,538
Spatially resolved transcriptomic profiling of degraded [mirzazadeh2023spatially] Bowel Visium 35 17,943
@@
A spatially resolved atlas of the human lung characterizes a gland-associated immune niche[madissoon2023spatially] Lung Visium 20 17,922
Molecular cartography uncovers evolutionary and microenvironmental dynamics in sporadic colorectal tumors[heiser_molecular_2023] Colon Visium 41 19,366
Spatially resolved multiomics of human cardiac niches[kanemaru_spatially_2023] Heart Visium 41 33,538
Transcriptome-scale spatial gene expression in the human dorsolateral prefrontal cortex[maynard2021transcriptome] Brain Visium 12 33,538
@@
Tertiary lymphoid structures generate and propagate anti-tumor antibody-producing plasma cells in renal cell cancer [meylan2022tertiary] Lymph node Visium 24 17,943
@@
Charting the Heterogeneity of Colorectal Cancer Consensus Molecular Subtypes using Spatial Transcriptomics: datasets [valdeolivas2023charting] Bowel Visium 14 36,601
Demo 10x Visium dataset for STQ [zenodo2024demo] Skin Visium 1 68,886
Nextflow Pipeline for Visium and H&E Data from Patient-Derived Xenograft Samples [Domanskyi2023.07.27.550727] Skin Visium 4 68886
Spotiphy: generative modeling in single-cell spatial whole transcriptomics [zenodo2024spotiphy] Brain Visium 2 32,285
@@ C.2 HEST-Library
The HEST-Library helps transform unstructured spatial transcriptomics and histology data into a unified format. An overview of the HEST-Library is provided in Figure 4. An example of fiducial detection is presented in Figure 5.
Figure 4: Overview of HEST-Library functionalities. HEST was designed to transform legacy data scrapped in multiple public repositories, such as NCBI, into unified HEST objects that can easily be integrated into computational pipelines.
Figure 5: Fiducial detection and automatic alignment in Visium. Corner fiducials on 6.5$\times$6.5mm and 11mm×$\times$11mm Visium slides are automatically detected with a finetuned Yolov8 model. The spot coordinates are then derived if at least 3 of the 4 corner fiducials are detected. This process enables automatically estimating the pixel resolution.
@@ C.3 HEST-Benchmark
Gene selection, XGBoost Forest, and Ridge regression models: We learn a regression model that maps the patch embeddings of each encoder to its corresponding gene expression profile. The XGboost model uses 100 estimators, a 0.1 learning rate, a max depth of 3, 0.8 subsampling, gamma of 0.0, regression alpha of 0.0, and regression lambda of 1.0. Additional information can be found in the XGBoost API [^12] https://xgboost.readthedocs.io/en/stable/python/python_api.html. The Ridge regression uses a fixed $L2$ regularization coefficient $\lambda$ set to $100/MC$, where $M$ is the embedding dimension and $C=50$ is the number of targets trained with the Regularized Least-Squares Routine solver (sklearn implementation). Both regression models are trained to predict a panel constituted of the 50 most variable genes of each task. Specifically, for each task, we select the 50 most variable genes across all spots and samples after excluding the genes that have non-zero counts in less than 10% of the spots.
@@
Task 1 [janesick2023high] IDC Breast 4 4 Xenium
@@
Task 5 [valdeolivas2023charting] COAD Colon 2 4 Xenium
Task 6 [valdeolivas2023charting] READ Rectum 2 4 Visium
Task 7 [meylan2022tertiary] ccRCC Kidney 24 24 Visium
Task 8 LUAD Lung 2 2 Xenium
Task 9 [liu2022single] IDC Axillary lymph nodes 4 4 Visium
@@
Task 1: Prediction of expression in invasive ductal carcinoma (breast cancer, IDC). We used all publicly available Xenium samples available on 10x Genomics (“FFPE Human Breast using the Entire Sample Area”, 2 patients) and two samples published in  [janesick2023high] (TENX95, TENX99, NCBI783, NCBI785). All samples are FFPE sections imaged with the Xenium pipeline v1.
@@
Task 2: Prediction of expression in prostate adenocarcinoma (prostate cancer, PRAD). We used all 23 Visium samples (fresh frozen sections) from 2 patients published in [erickson2022spatially] (MEND139 to MEND162). Both patients were diagnosed with prostatic acinar adenocarcinoma with a (4+3) Gleason score (ISUP group 4).
@@
Task 3: Prediction of expression in pancreatic adenocarcinoma (pancreatic cancer, PAAD). We used 3 samples from 3 different patients from 10x Genomics (“FFPE Human Pancreas with Xenium Multimodal Cell Segmentation” and “Pancreatic Cancer with Xenium Human Multi-Tissue and Cancer Panel”). All samples are FFPE sections processed with Xenium pipeline v1 (TENX116, TENX126, TENX140).
@@
Task 4: Expression prediction in skin cutaneous melanoma (skin cancer, SKCM). We used 2 samples from 2 different patients from 10x Genomics website (“Human Skin Data with Xenium Human Multi-Tissue and Cancer Panel”). All samples are FFPE sections processed with Xenium pipeline v1 (TENX115, TENX117).
@@
Task 5: Prediction of expression in colon adenocarcinoma (colon cancer, COAD). We used 4 COAD samples from 2 different patients available on 10x Genomics (TENX111, TENX147, TENX148, TENX149). All samples are fresh frozen sections processed with Visium.
@@
Task 6: Prediction of expression in rectal adenocarcinoma (rectum cancer, READ). We used 4 READ samples from 2 different patients published in [valdeolivas2023charting]. All samples are fresh frozen sections processed with Visium (ZEN36, ZEN40, ZEN48, ZEN49).
@@
Task 7: Prediction of expression in clear cell renal cell carcinoma (kidney cancer, ccRCC). We used the 24 ccRCC samples of 24 different patients published in [meylan2022tertiary]. All samples are fresh frozen sections processed with Visium (INT1 to INT24).
@@
Task 8: Prediction of expression in lung adenocarcinoma (lung cancer, LUAD). We used 2 LUAD samples from 2 different patients from 10x genomics (“Preview Data: FFPE Human Lung Cancer with Xenium Multimodal Cell Segmentation”). All samples are fresh frozen sections processed with Xenium pipeline v1 (TENX118, TENX141).
@@
Task 9: Prediction of expression in axillary lymph nodes in IDC patients. We used 4 axillary lymph node samples from 2 IDC patients published in [liu2022single]. All samples are fresh frozen sections processed with Visium (NCBI681, NCBI682, NCBI683, NCBI684).
@@
Table A12: State-of-the-art foundation models for histology evaluated on HEST-Benchmark. B: Base, L: Large, H: Huge, G: Giant. : number of patches during pretraining, ∗∗: number of patches during fine-tuning.
Name Number of Number of Magnification Model Training Number of
slides patches recipe parameters
ResNet50 (IN) [lu2021data] N/A 1.2M N/A ResNet-50 Supervised 23M
CTransPath [wang2021transpath] 32k 17M 10$\times$ Swin-T MoCov3 28M
Remedis [azizi2023robust] 10k 10M 20$\times$ ResNet-152 iBOT 232M
Phikon [filiot2023scaling] 6k 43.3M 20$\times$ ViT-B iBOT 86M
UNI [chen2024towards] 100k 100M 20$\times$ ViT-L DINOv2 307M
CONCH [lu2024towards] N/A 16M + 1.17M∗∗ Many ViT-B CoCa 86M
GigaPath [xu2024wholeslide] 171k 1.3B 20$\times$ ViT-g DINOv2 1.13B
Virchow [vorontsov2024foundation] 1.5M 2B 20$\times$ ViT-H DINOv2 632M
Virchow 2 [Zimmermann2024Virchow2SS] 3.1M 1.9B Many ViT-H DINOv2 632M
H-Optimus-0 500K 273M 20x ViT-g DINOv2 1.13B
UNIv1.5 350K 432M 20x ViT-g DINOv2 1.13B
@@
We provide a brief description of each patch encoder assessed with the HEST-Benchmark.
@@
ResNet50 (IN) [lu2020data]: This model uses a ResNet50 backbone [he2016deep] trained on ImageNet [deng2009imagenet] (1.2 million natural images). Following prior work [lu2020data], the patch embeddings are extracted by taking the representation at the penultimate layer before final classification.
@@
CTransPath [wang2022transformer]: This model uses a “Tiny” Swin Transformer backbone [liu2021swin] with a window size of 14 (Swin-T/14, 28 million parameters) pretrained on TCGA and PAIP datasets (17 million images) using MoCoV3 [chen2021an].
@@
Remedis [azizi2023robust]: This model uses a ResNet-152$\times$2 (232 million parameters) initialized with the “Big Transfer”-medium protocol [kolesnikov2020big] on ImageNet-22K and pretrained with SimCLR [chen2020simple] on TCGA.
@@
Phikon [filiot2023scaling]: This model uses a Vision Transformer-Base (ViT-B, 86 million parameters) [dosovitskiy2021image] trained on TCGA data using iBOT [zhou2021ibot].
@@
UNI [chen2024towards]: This model uses a ViT-Large (ViT-L, 307 million parameters) [dosovitskiy2021image] trained on 100 million histology images (over 100,000 slides) from proprietary and public data using DINOv2 [oquab2023dinov2].
@@
CONCH [lu2024towards]: This model uses a ViT-B (86 million parameters) trained on a smaller version of UNI using iBOT, and then fine-tuned on 1.17 million histology image–caption pairs extracted from online educational and research resources using CoCa [yu2022coca].
@@
GigaPath [xu2024wholeslide]: This model uses a ViT-giant (1.13 billion parameters) trained on 1.3 billion image patches from 171,189 WSIs at 20$\times$ magnification using DINOv2.
Virchow [vorontsov2024foundation]: This model uses a ViT-Huge (632M parameters) trained on 2 billion image patches and 1.5M WSIs at 20$\times$ magnification using DINOv2.
Virchow 2 [Zimmermann2024Virchow2SS]: This model uses a ViT-Huge (632M parameters) trained on 1.9B patches and 3.1M WSIs using DINOv2
H-Optimus-0: This model uses a ViT-giant (1.13B parameters) trained 273 million image patches from 500,000 WSIs at 20$\times$ magnification using DINOv2.
UNIv1.5: This model uses a ViT-giant (1.13B parameters) trained on 432 million image patches from 350,000 WSIs using DINOv2.
@@ Appendix D HEST for multimodal representation learning
We provide additional information regarding CONCH fine-tuning using multimodal alignment. CONCH-FT model, a ViT-Base model initialized with CONCH weights, was fine-tuned for 50 epochs using a cosine learning rate scheduler, with a base learning rate of $10^{-4}$ for the image encoder and $10^{-3}$ for the expression encoder. Only the last 3 layers of the model were fine-tuned, with a layer-wise learning decay rate of 0.7. For training with the infoNCE loss, a contrastive temperature of $10^{-2}$ and batch size of 1,024 pairs of patch and transcriptomics were used. A combination of random horizontal/vertical flip and color jittering was employed for image patch augmentation.
The rank of the embedding space (also referred to as smooth rank measure [garrido2023rankme]) measures the quality of the embeddings produced from encoders trained in unsupervised or self-supervised manners. Given the patch embedding matrix $H\in\mathcal{R}^{N\times d}$ and $d<N$, where $N$ is the number of patches and $d$ is the feature dimension, we compute the rank as the entropy of the $d$ L1-normalized singular values of $H$.
@@ Appendix E HEST for discovery
Cells were segmented and classified using CellViT [horst2024cellvit]. To find the gene expression profile of each neoplastic cell, we matched each cell to its corresponding cell index in Xenium by assigning the index for which the distance between the cell centroids was the smallest. After matching all neoplastic cells, only those cells for which the assignment was unique were kept. After this filtering step, an average of 91% of the cells per sample were kept while 9% of the cells were discarded.
@@ F.1 Motivation for dataset creation
  • Why was the dataset created? HEST-1k was designed with three key applications: (1) multimodal representation learning of histology and transcriptomics, (2) biomarker exploration and characterization, and (3) benchmarking foundation models for pathology. Despite many publicly available resources, no existing unified and user-friendly formatting was available to bring ST into the world of deep learning.
  • What (other) tasks could the dataset be used for? Are there obvious tasks for which it should not be used? Users are welcome to introduce new, creative ways to use the dataset. However, users are not allowed to try to retrieve patient information from the existing data. A dedicated section is provided to discuss ethical considerations and intended usage.
  • Has the dataset been used for any tasks already? If so, where are the results so others can compare (e.g., links to published papers)? The metadata attached to HEST-1k reports all samples that were made public as part of a publication (peer-reviewed or not).
  • Who funded the creation of the dataset? HEST is supported by the Brigham and Women’s Hospital (BWH) President’s Fund, Mass General Hospital (MGH) Pathology, and the National Institute of Health (NIH) National Institute of General Medical Sciences (NIGMS) through R35GM138216.